# VeriThoughts: Enabling Automated Verilog Code Generation using Reasoning and Formal Verification

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in generating software code (such as in programming languages like Python, Java, and C++) from natural language prompts 1. This has resulted in widespread adoption of these tools in real-world software engineering workflows, improving developer productivity by up to 55.855.855.855.82. However, notwithstanding a growing body of research, code generation for hardware design languages (HDL) has proven to be much more challenging. Verilog, one of the most commonly used HDLs, allows designers to specify the function of a chip at a high level, leaving its conversion to a manufacturable circuit to automated tools. Even so, HDL programming is notoriously tedious and time-consuming 3. LLM-based automatic Verilog code generation would therefore significantly improve chip design productivity.
A key challenge in automating HDL code generation is the scarcity of HDL codes on the web compared to languages like Python 4. Code generation models trained on open-source code repositories tend to struggle on Verilog problems 5. Recent efforts have sought to build both training and evaluation datasets to remedy this issue 5, 6, 4, 7, 8. An early effort, VeriGen 4, scraped more than 108K Verilog files from GitHub, which were then used to fine-tune open-source LLMs using self-supervised next token prediction. A small evaluation dataset of 17 natural language prompts, along with test inputs and desired outputs was proposed to benchmark the accuracy of fine-tuned LLMs. Larger training sets of HDL/Verilog code 7 and smaller evaluation datasets of prompts and corresponding test inputs/outputs 9, 8 have been released.
Training sets containing Verilog code alone 6, 10 can only be used to finetune LLMs with self-supervised next token prediction, but not for subsequent supervised fine-tuning (SFT) steps that are important for performance 11. For SFT, pairs of natural language prompts (or Verilog coding questions) and corresponding Verilog implementations are needed. Manually writing descriptive prompts/questions for thousands of Verilog modules is infeasible, and automated prompt/question generation methods have been plagued by hallucinations 12, 13, 14 The recent success of reasoning models like DeepSeek-R1 15 and Qwen3 16 suggests that SFT on chain-of-thought (CoT) traces can improve performance, but CoT traces for Verilog code generation are not publicly available. Moreover, evaluation datasets require mechanisms to check the correctness of generated code—this has typically been done using test inputs/outputs in both the software and hardware communities. But test inputs/outputs may not be readily available for code scraped from open repositories, and LLM generated tests are known to be buggy 17.
@@
Our data generation process shown in Figure 1 starts with MetRex 10, one of the largest datasets of “synthesizable" Verilog (i.e., Verilog that can be mapped to a hardware implementation) collated from many sources. Like VeriGen, Metrex has Verilog code only. Using a frontier Gemini model, we first produce a question ($Q$) to which a MetRex code sample ($V$) is a correct response. A second model, Deepseek-R1-670B, is used to generate a reasoning trace ($R$) and candidate Verilog implementation ($V$) responsive to the question ($Q)$. However, the question and/or the generated Verilog code might be incorrect due to hallucinations, resulting in a noisy dataset. To address this concern, we propose a formal verification-based “self-consistency" check that compares ground-truth Verilog with generated Verilog; a mismatch indicates that one or more among $Q$, $R$ or $V^{*}$ are incorrect. Manual analysis of a random subset of training data confirms value of these labels. We demonstrate the value of our dataset by enabling answers to interesting RQs about the usefulness of reasoning traces and self-consistency labels, and by training SoTA open-source models for Verilog generation.
@@ 2 Background & Related Work
LLMs have demonstrated strong capabilities in code generation for programming languages such as C and Python 18, 19, 20, 21, 22, 23, 24. This success is largely attributed to the availability of extensive training datasets comprising source code in one or multiple programming languages. These datasets can span several hundreds of gigabytes, enabling LLMs to learn syntactic patterns, semantic structures, and common usage conventions. Inputs to the models include natural language instructions, comments, partial code snippets, or combinations thereof.
Researchers explored fine-tuning open-source LLMs for HDL generation 25, 12, 26. To assess performance, benchmarks such as Verilog Eval 13 have been introduced to evaluate Verilog generation quality. Despite these efforts, even fine-tuned models often lag behind leading closed-source, proprietary LLMs such as Claude Sonnet or the GPT series in output quality and reliability.
In order to enable training of domain-specific Verilog generation models, a number of relevant datasets have been created. The Verilog GitHub dataset 25 includes approximately 108,000 Verilog files scraped from GitHub, and has been used to train models such as VeriGen and CL-Verilog. Building on this dataset, MG-Verilog introduced a multi-grained version of the Verilog GitHub dataset with corrected syntax. The RTL-Coder Resyn dataset utilizes LLMs to create a number of machine-generated Verilog samples 12. Finally, the MetRex dataset, which we build upon here, contains over 25,000 Verilog files which are synthesizable via Yosys, and includes natural language descriptions of all relevant post-synthesis metrics. Table 1 contains comparisons across existing large-scale Verilog datasets, including the number of Verilog designs they include, whether they have been checked for synthesizability, whether they include natural language descriptions, whether they include a reasoning trace related to each design, and whether they include self-consistency checks.
Table 1: Comparison across existing large-scale Verilog datasets. Factors include synthesizability checking, natural language questions/descriptions, availability of reasoning traces, and self-consistency checking.
Dataset Design Count Synth. Verilog Nat. Lang. Qns. Reasoning Self-Cons. Check
Verilog GitHub (VeriGen) 108 971108971108\,971108 971
MG-Verilog 11 1441114411\,14411 144
Resyn (RTL-Coder) 26 5322653226\,53226 532
MetRex 25 8682586825\,86825 868
VeriThoughts (train) 20 1732017320\,17320 173
@@ 3 VeriThoughts Dataset
We now introduce VeriThoughts, a unique, large-scale, formally-verified Verilog reasoning dataset. VeriThoughts is comprised of 20K samples of Verilog RTL code, each paired with a prompt describing the code, newly generated Verilog code from the prompt, the reasoning traces used to generate the new Verilog code, and a label indicating whether the generated Verilog and original Verilog are functionally equivalent. The original RTL code, $V$, is sourced from the MetRex dataset 10. We then generate prompts, $Q$, for each Verilog entry with the Gemini-2.0-Flash-Thinking-Experimental model. These prompts are then given to Deepseek-R1 which generates Verilog code $V^{*}$ and reasoning trace $R$. Finally, we perform formal verification $E$ of the original and generated Verilog code with the Yosys framework to obtain a self-consistency label $L_{C}\leftarrow E(V=V^{*})$. An overview of this framework can be seen in Figure 1.
@@ 3.1 MetRex: Original Verilog Dataset
The MetRex dataset, released under the BSD 3-Clause license 10, contains 25.8K Verilog designs taken from publicly available sources. These include machine generated designs such as the RTL-Coder dataset 7 and human created designs such as the VeriGen dataset 4. The dataset creators took scraped Verilog designs from the web and cleaned the data to ensure that only synthesizable designs remained. Note that some of the RTL included in this dataset 7 include LLM-generated descriptions of their function. However, the quality of these annotations is unverified and for our purposes, we only use the RTL, and discard all existing generated text descriptions. We filter the dataset to around 25K entries that all have a character length of less than 10,000 characters.
@@ 3.2 Generating Prompts for Unlabeled Verilog Code
Notwithstanding incorrect questions, other work on software code generation has found that even inconsistent question-answer pairs in a training dataset can help improve code generation accuracy 27. To help resolve the issue in the context of Verilog code generation, VeriThoughts also includes additional labels that attempt to capture the impact of consistency between pairs of $(V,Q)$. Unfortunately, manual annotations are not feasible for a dataset of our size—instead, we employ self-consistency check, described below, as a proxy for question answer consistency.
@@ 3.6 Manual Analysis of VeriThoughts Subset
We provide three representative examples to illustrate the prompt error (Fig. 4), code-gen error (Fig. 5), and both errors (Fig. 6) cases identified above. Among the $50$ mismatched data points, most of them ($66\%$) correspond to the prompt error case: the prompt produced by the LLM cannot accurately explain the golden Verilog code. A smaller fraction ($16$) correspond to the code-generation error case, where the prompt is accurate but the LLM implements it inexactly. Finally, $18$ belong to the both errors case; here a common pattern is that the generated prompt omits some implementation details of the golden Verilog, and the LLM fills in the missing details, leading to function mismatches.
@@ 4.1 Experimental Setup
In all subsequent experiments we perform supervised fine-tuning (SFT) of Qwen-2.5-Instruct-7B models using a learning rate of $8*10^{-5}$, a cosine scheduler, bf16 format, and three epochs over our training set. This finetuning is done with Llama-Factory 28. The full fine-tuning was performed on 8xA100 GPUs and takes roughly 2-4 hours for instruction tuned models and roughly 7-14 hours for reasoning models.
@@
We evaluate both preexisting and our trained models on two benchmarks. VerilogEval 5 has 156 problems sourced from HDLBits and is meant to test a language model’s ability to generate a diverse distribution of Verilog code. We focus on the "Human" split of VerilogEval 1.0 because it has higher-quality labels and problems than the machine-generated "Machine" split. The benchmark is evaluated using human-designed test cases for every Verilog problem.
In addition to VerilogEval, we create a new benchmark sourced by randomly sampling a hold-out set of 291 questions from VeriThoughts. We call this the VeriThoughts benchmark, and validate it using the formal verification scheme described in Section 3. This helps avoid the strain of manually creating unit test cases (such as in VerilogEval) while also providing a stronger level of verification for our benchmark. We generate responses for VerilogEval with a temperature of 0.5, top_p of 0.90, and maximum generation length of 1024 tokens. This is in line with the VerilogEval Human 1.0 results 13. For our VeriThoughts benchmark, we generate responses with a temperature of 0.6, top_p of 0.95, and maximum generation length of 16384. Temperature and top_p are in line with the suggested generation settings for DeepSeek and the maximum generation length is increased for longer problems (and associated reasoning traces) in VeriThoughts. We report standard "pass@k" 29 metrics for $k=\{1,5,10\}$, standard values used in the literature. Pass@k is evaluated over 20 trials.
@@ 4.2 Research Questions
Reasoning in LLMs has been shown to help in tasks such as coding and science 30, 31; does this apply to a low-resource coding language such as Verilog? We validate this hypothesis by comparing models trained on full reasoning traces versus instruction tuning along. The results are in Table 4.
@@ 4.3 A new state of the art on VerilogEval
Qwen3-14B Reasoning 87.4% 96.8% 98.1% 28.1% 45.6% 51%
VeriThoughts-14B (Ours) Reasoning 78.5% 90.0% 92.1% 43.7% 52.2% 55.14%

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in generating software code (such as in programming languages like Python, Java, and C++) from natural language prompts Hou24. This has resulted in widespread adoption of these tools in real-world software engineering workflows, improving developer productivity by up to 55.855.8Peng23. However, notwithstanding a growing body of research, code generation for hardware design languages (HDL) has proven to be much more challenging. Verilog, one of the most commonly used HDLs, allows designers to specify the function of a chip at a high level, leaving its conversion to a manufacturable circuit to automated tools. Even so, HDL programming is notoriously tedious and time-consuming joshi1995measuring. LLM-based automatic Verilog code generation would therefore significantly improve chip design productivity.
A key challenge in automating HDL code generation is the scarcity of HDL codes on the web compared to languages like Python Thakur24. Code generation models trained on open-source code repositories tend to struggle on Verilog problems Liu23. Recent efforts have sought to build both training and evaluation datasets to remedy this issue Liu23, Thakur23, Thakur24, Liu24, Pinckney25. An early effort, VeriGen Thakur24, scraped more than 108K Verilog files from GitHub, which were then used to fine-tune open-source LLMs using self-supervised next token prediction. A small evaluation dataset of 17 natural language prompts, along with test inputs and desired outputs was proposed to benchmark the accuracy of fine-tuned LLMs. Larger training sets of HDL/Verilog code Liu24 and smaller evaluation datasets of prompts and corresponding test inputs/outputs Lu24, Pinckney25 have been released.
Training sets containing Verilog code alone Thakur23, Abdelatty25 can only be used to finetune LLMs with self-supervised next token prediction, but not for subsequent supervised fine-tuning (SFT) steps that are important for performance Yang24. For SFT, pairs of natural language prompts (or Verilog coding questions) and corresponding Verilog implementations are needed. Manually writing descriptive prompts/questions for thousands of Verilog modules is infeasible, and automated prompt/question generation methods have been plagued by hallucinations liu2024rtlcoder, liu2023verilogeval, zhang2024mgverilogmultigraineddatasetenhanced The recent success of reasoning models like DeepSeek-R1 DeepSeekR1 and Qwen3 qwen3 suggests that SFT on chain-of-thought (CoT) traces can improve performance, but CoT traces for Verilog code generation are not publicly available. Moreover, evaluation datasets require mechanisms to check the correctness of generated code—this has typically been done using test inputs/outputs in both the software and hardware communities. But test inputs/outputs may not be readily available for code scraped from open repositories, and LLM generated tests are known to be buggy Liu2024LLMPoweredTC.
@@
Our data generation process shown in Figure 1 starts with MetRex Abdelatty25, one of the largest datasets of “synthesizable" Verilog (i.e., Verilog that can be mapped to a hardware implementation) collated from many sources. Like VeriGen, Metrex has Verilog code only. Using a frontier Gemini model, we first produce a question ($Q$) to which a MetRex code sample ($V$) is a correct response. A second model, Deepseek-R1-670B, is used to generate a reasoning trace ($R$) and candidate Verilog implementation ($V$) responsive to the question ($Q)$. However, the question and/or the generated Verilog code might be incorrect due to hallucinations, resulting in a noisy dataset. To address this concern, we propose a formal verification-based “self-consistency" check that compares ground-truth Verilog with generated Verilog; a mismatch indicates that one or more among $Q$, $R$ or $V^{*}$ are incorrect. Manual analysis of a random subset of training data confirms value of these labels. We demonstrate the value of our dataset by enabling answers to interesting RQs about the usefulness of reasoning traces and self-consistency labels, and by training SoTA open-source models for Verilog generation.
@@ 2 Background & Related Work
LLMs have demonstrated strong capabilities in code generation for programming languages such as C and Python chen2021evaluating, nijkamp2022codegen, feng2020codebert, fried2022incoder, wang2021codet5, manyika2023overview, roziere2023code. This success is largely attributed to the availability of extensive training datasets comprising source code in one or multiple programming languages. These datasets can span several hundreds of gigabytes, enabling LLMs to learn syntactic patterns, semantic structures, and common usage conventions. Inputs to the models include natural language instructions, comments, partial code snippets, or combinations thereof.
Researchers explored fine-tuning open-source LLMs for HDL generation thakur2023verigen, liu2024rtlcoder, nakkab2024romebuiltsinglestep. To assess performance, benchmarks such as Verilog Eval liu2023verilogeval have been introduced to evaluate Verilog generation quality. Despite these efforts, even fine-tuned models often lag behind leading closed-source, proprietary LLMs such as Claude Sonnet or the GPT series in output quality and reliability.
In order to enable training of domain-specific Verilog generation models, a number of relevant datasets have been created. The Verilog GitHub dataset thakur2023verigen includes approximately 108,000 Verilog files scraped from GitHub, and has been used to train models such as VeriGen and CL-Verilog. Building on this dataset, MG-Verilog introduced a multi-grained version of the Verilog GitHub dataset with corrected syntax. The RTL-Coder Resyn dataset utilizes LLMs to create a number of machine-generated Verilog samples liu2024rtlcoder. Finally, the MetRex dataset, which we build upon here, contains over 25,000 Verilog files which are synthesizable via Yosys, and includes natural language descriptions of all relevant post-synthesis metrics. Table 1 contains comparisons across existing large-scale Verilog datasets, including the number of Verilog designs they include, whether they have been checked for synthesizability, whether they include natural language descriptions, whether they include a reasoning trace related to each design, and whether they include self-consistency checks.
Table 1: Comparison across existing large‐scale Verilog datasets. Factors include synthesizability checking, natural language questions/descriptions, availability of reasoning traces, and self-consistency checking.
Dataset
Design
Count
Synth.
Verilog
Nat. Lang.
Qns.
Reasoning
Self-Cons.
Check
Verilog GitHub (VeriGen) 108 971108\,971
MG-Verilog 11 14411\,144
Resyn (RTL-Coder) 26 53226\,532
MetRex 25 86825\,868
VeriThoughts (train) 20 17320\,173
@@ 3 VeriThoughts Dataset
We now introduce VeriThoughts, a unique, large-scale, formally-verified Verilog reasoning dataset. VeriThoughts is comprised of 20K samples of Verilog RTL code, each paired with a prompt describing the code, newly generated Verilog code from the prompt, the reasoning traces used to generate the new Verilog code, and a label indicating whether the generated Verilog and original Verilog are functionally equivalent. The original RTL code, $V$, is sourced from the MetRex dataset Abdelatty25. We then generate prompts, $Q$, for each Verilog entry with the Gemini-2.0-Flash-Thinking-Experimental model. These prompts are then given to Deepseek-R1 which generates Verilog code $V^{*}$ and reasoning trace $R$. Finally, we perform formal verification $E$ of the original and generated Verilog code with the Yosys framework to obtain a self-consistency label $L_{C}\leftarrow E(V=V^{*})$. An overview of this framework can be seen in Figure 1.
@@ 3.1 MetRex: Original Verilog Dataset
The MetRex dataset, released under the BSD 3-Clause license Abdelatty25, contains 25.8K Verilog designs taken from publicly available sources. These include machine generated designs such as the RTL-Coder dataset Liu24 and human created designs such as the VeriGen dataset Thakur24. The dataset creators took scraped Verilog designs from the web and cleaned the data to ensure that only synthesizable designs remained. Note that some of the RTL included in this dataset Liu24 include LLM-generated descriptions of their function. However, the quality of these annotations is unverified and for our purposes, we only use the RTL, and discard all existing generated text descriptions. We filter the dataset to around 25K entries that all have a character length of less than 10,000 characters.
@@ 3.2 Generating Prompts for Unlabeled Verilog Code
Notwithstanding incorrect questions, other work on software code generation has found that even inconsistent question-answer pairs in a training dataset can help improve code generation accuracy muennighoff2025s1simpletesttimescaling. To help resolve the issue in the context of Verilog code generation, VeriThoughts also includes additional labels that attempt to capture the impact of consistency between pairs of $(V,Q)$. Unfortunately, manual annotations are not feasible for a dataset of our size—instead, we employ self-consistency check, described below, as a proxy for question answer consistency.
@@ 3.6 Manual Analysis of VeriThoughts Subset
We provide three representative examples to illustrate the prompt error (Fig. 4), code-gen error (Fig. 5), and both errors (Fig. 6) cases identified above. Among the $50$ mismatched data points, most of them ($66\%$) correspond to the prompt error case: the prompt produced by the LLM cannot accurately explain the golden Verilog code. A smaller fraction ($16$ %) correspond to the code-generation error case, where the prompt is accurate but the LLM implements it inexactly. Finally, $18$ % belong to the both errors case; here a common pattern is that the generated prompt omits some implementation details of the golden Verilog, and the LLM fills in the missing details, leading to function mismatches.
@@ 4.1 Experimental Setup
In all subsequent experiments we perform supervised fine-tuning (SFT) of Qwen-2.5-Instruct-7B models using a learning rate of $8*10^{-5}$, a cosine scheduler, bf16 format, and three epochs over our training set. This finetuning is done with Llama-Factory zheng2024llamafactory. The full fine-tuning was performed on 8xA100 GPUs and takes roughly 2-4 hours for instruction tuned models and roughly 7-14 hours for reasoning models.
@@
We evaluate both preexisting and our trained models on two benchmarks. VerilogEval Liu23 has 156 problems sourced from HDLBits and is meant to test a language model’s ability to generate a diverse distribution of Verilog code. We focus on the "Human" split of VerilogEval 1.0 because it has higher-quality labels and problems than the machine-generated "Machine" split. The benchmark is evaluated using human-designed test cases for every Verilog problem.
In addition to VerilogEval, we create a new benchmark sourced by randomly sampling a hold-out set of 291 questions from VeriThoughts. We call this the VeriThoughts benchmark, and validate it using the formal verification scheme described in Section 3. This helps avoid the strain of manually creating unit test cases (such as in VerilogEval) while also providing a stronger level of verification for our benchmark. We generate responses for VerilogEval with a temperature of 0.5, top_p of 0.90, and maximum generation length of 1024 tokens. This is in line with the VerilogEval Human 1.0 results liu2023verilogeval. For our VeriThoughts benchmark, we generate responses with a temperature of 0.6, top_p of 0.95, and maximum generation length of 16384. Temperature and top_p are in line with the suggested generation settings for DeepSeek and the maximum generation length is increased for longer problems (and associated reasoning traces) in VeriThoughts. We report standard "pass@k" chen2021codex metrics for $k=\{1,5,10\}$, standard values used in the literature. Pass@k is evaluated over 20 trials.
@@ 4.2 Research Questions
Reasoning in LLMs has been shown to help in tasks such as coding and science sky_t1_2025, Open_Thoughts; does this apply to a low-resource coding language such as Verilog? We validate this hypothesis by comparing models trained on full reasoning traces versus instruction tuning along. The results are in Table 4.
@@ 4.3 A new state of the art on VerilogEval
Qwen3-14B Reasoning 87.4% 96.8% 98.1% 28.1% 45.6% 51.0%
VeriThoughts-14B (Qwen2.5 Base) Reasoning 78.5% 90.0% 92.1% 43.7% 52.2% 55.14%
@@ Appendix A Appendix
Table 8: Pass@k scores for closed- and open-source models on Verilog Eval. Closed source models first, followed by open source models.
Model Model Type Verilog Eval
Pass@1 Pass@5 Pass@10
Claude 3.7 Sonnet Reasoning 66.3% 75.8% 78.5%
GPT-4o Instruct (MoE) 56.3% 66.7% 69.0%
GPT-4 Instruct (MoE) 43.5% 55.8% 58.9%
VeriThoughts-32B (Qwen2.5 Base) Reasoning 52.0% 61.6% 63.6%
VeriThoughts-32B (Qwen3 Base) Reasoning 49.5% 59.2% 61.9%
VeriThoughts-14B (Qwen3 Base) Reasoning 38.7% 51.8% 55.6%
RTLCoder-v1.1-7B Instruct 34.6% 43.4% 45.5%
RTLCoder-DeepSeek-v1.1-7B Instruct 39.7% 49.3% 51.9%